Solve the equation m2 +n2=29 and m+n =7

1
Expert's answer
2022-01-31T16:21:43-0500

Solve the equations:

m2 + n2 = 29

m + n = 7


First equation we could form in and simplified like this:

(m2 + n2 + 2mn) - 2mn= 29

(m + n)2 - 2mn = 29


We know that m + n = 7, so then (m + n)2 - 2mn = 29 becomes:

72 - 2mn = 29

49 - 2mn = 29

2mn = 20

m*n = 10


Now we have :

m * n = 10

m + n = 7


And the numbers for m and n could only be 2 and 5 or 5 and 2.


Answer: (2, 5), (5, 2).